oC b) 75 C- 200°C d) 200°F oC V. Calculate the percent error for the following situations. a) During an experiment a scientist finds the density of a rock to be 3.29 g/em2. The accepted value for this type of rock is 3.39 g/cm. Calculate the percent error A brick has a mass of 500 grams written on it. When the mass is measured it is found to be 489 grams. Calculate the percent error A student performing an experiment about the melting point of an object. After 5 trials he found the melting point t be 192 °C, 206 °C, 197 °C, 199 °C, and 202 °C. The accepted melting value temperature is 200 C. Calculate the percent error. b) For each pair of units circle or highlight the smallest quantity. a) decimeter and dekameter b) centimeter and millimeter
Answer #1

(a) 75 deg F = (75-32)*5/9 = 23.9 deg C

(b) 75 deg C = 75 * (9/5) +32 = 167 deg F

(c) 200 deg C = 200 * (9/5) +32 = 392 deg F

(d) 200 deg F = (200-32)*5/9 = 93.3 deg C

IV (a) Percentage error = [(Accepted value - measure value) / Accepted value] x 100

= [(3.39 - 3.29) / 3.39] x 100 = 2.95 %

(b) Like above, percentage error = [(500 - 489)/500] x 100 = 2.20 %

(c) Average measured value = (192 + 206 + 197 + 199 + 202)/ 5 = 199.2 deg C

Accepted melting point = 200 deg C

So, percentage error = [(200 - 199.2) / 200]x100 = 0.40 %

(V) (a) Smallest quantity - decimeter

(b) millimeter.